how do you solve |2a+7|=a-4
(please provide answer also)

Answer by stanbon(75887)

|2a+7|=a-4
----------
2a+7 = a-4 or 2a+7 = 4-a
a = -11 or 3a = -3
a = -11 or a = -1
---------------------
Checking:
a = -11 does not work in the original equation.
a = -1 does not work in the original equation.
---
Conclusion:
No solution.
=================
Cheers,
Stan H.